# A Practical Synthesis of Detecting AI-Generated Textual, Visual, and Audio Content

### v1

@@ III-B1 LLM prompting and zero-shot methods
One popular approach leverages an external LLM to classify whether a piece of text is AI-generated or not [12]. With carefully crafted prompts, these zero-shot methods can achieve moderate accuracy quickly. However, sensitivity to prompt design and the adversarial gap between the LLM detector and the text generator are common issues [22].
@@ III-B4 Watermarking for AI text
Cooperative watermarking modifies text generation at token selection time, embedding an imperceptible pattern in the distribution of words or punctuation [30, 31]. A verifier can detect such patterns after the fact. While promising for major industrial LLMs that adopt the standard, watermarking fails if malicious or open-source models do not embed it, or if paraphrasing disrupts the signal [32, 33].
@@ IV-B4 Watermarking for visual media
When the generative pipeline is compliant, watermarks are embedded. These vary from invisible spatial pixel encodings to frequency manipulations [30, 40]. However, simple image transformations, such as crop or rotation, can weaken naive watermarks [32, 41] unless specifically designed for robustness.
IV-B5 Video deepfake detection
Beyond static artifacts in frames, deepfake videos exhibit temporal inconsistencies in facial expressions, motion, or geometry [42, 43, 44]. Optical flow analysis or 3D geometry can spot unnatural transitions [45, 46]. Some methods track facial micro-expressions (e.g., eye blinking [47]), aligning them with known patterns of real human physiology.
@@ IV-C Datasets and Benchmarks
Well-known datasets include CelebA-HQ [48] for facial images and LAION [49] for broad image domains. Large-scale synthetic benchmarks like StyleGAN or Stable Diffusion outputs help train classifiers [17]. For video, FaceForensics++ [50], DFDC [51], and Celeb-DF are commonly used [44]. However, frequent advancement of the generative model requires continuous expansion of the detection dataset.
@@ V-B1 Pipeline classifiers
Features such as Mel frequency cepstral coefficients (MFCC), Linear frequency cepstral coefficients (LFCC) spectrogram-based descriptors are extextracted and then fed into machine learning models (SVM, XGBoost, or CNN) [14, 8]. They typically rely on analyzing subtle cues in pitch, timbre, and vocal fold dynamics.
@@ V-B4 Audio watermarking
V-C Robustness and Challenges
Generalization across accents and languages is a key challenge, as a universal voice deepfake detector must handle wide linguistic and accent diversity [61]. Another major constraint is time constraints and real-time detection, since fraud scenarios may require near-instant detection in phone calls or chat streams, while many advanced models remain too slow for real-time deployment. Furthermore, music complexity adds another layer of difficulty—music contains higher dimensionality involving harmonies, instrumentation, and lyrics, which demands specialized features and large labeled datasets [16].
@@ VII-A Unified Themes Across Modalities
Despite modality differences, several consistent themes arise. (1) Arms race with generators: As generative models advance, older detection strategies degrade, requiring frequent retraining or adaptation [4, 63]. (2) Vulnerability to perturbations: Slight paraphrasing in text, minor image edits (e.g., crop and color shift), or audio pitch/time adjustments can bypass naive methods [21, 38]. (3) Watermarking as partial solution: Watermarking is useful if widely adopted by model providers, but it’s easily circumvented by uncooperative or malicious providers [33]. (4) Ensemble or multimodal methods: Combining multiple cues (statistical, watermark, contextual) and bridging text–image–audio modalities yields more robust detection [34, 35]. (5) Ethical pitfalls: Systemic large-scale scanning can infringe on user privacy, and erroneous misclassifications can harm reputations.
VII-B Emerging Approaches: Self-Supervised Learning and Generative Model Fingerprinting
Recent research has begun exploring self-supervised learning techniques to enhance detection robustness. Approaches leveraging pretext tasks (e.g., contrastive learning) enable detectors to learn robust feature representations from unlabeled data, capturing subtle discrepancies in AI-generated content. Similarly, generative model fingerprinting techniques aim to extract unique signatures inherent to specific generator architectures or training pipelines. These methods hold promise for improved performance, particularly in cross-domain and evolving generative scenarios.
VII-C Open Problems and Research Gaps
VII-C1 Adversarially robustness
Most detection tools remain sensitive to small input perturbations or domain shifts. Research on adversarial training or specialized defenses in text or multimedia deepfake detection is nascent [29, 38].
VII-C2 Domain generalization
Detectors generally excel on known generator architectures or stable distributions. But in real deployment, content can come from newly released models. Future solutions may rely on generative model fingerprinting or dynamic unsupervised adaptation [64].
VII-C3 Explainability and calibration
High-stakes contexts (legal, academic) demand interpretable detection. Visualizing suspicious words, image regions, or spectral frames fosters trust. Score calibration ensures confidence estimates are meaningful, limiting false accusations.
VII-C4 Lack of diversity in benchmark datasets
For text, English is overrepresented; for images, faces dominate available data. Audio datasets often focus on one or two languages. This scarcity undermines generalization to real global use cases [3, 8].
VII-C5 Regulatory and policy dimensions
Debates abound regarding mandated watermarking or content labeling, standardizing detection APIs, or clarifying liability for generative misuse. Government bodies and industry coalitions may adopt guidelines influencing detection approach design.
VII-D Future Directions
Future research in AI content detection is exploring several promising directions. (1) Retrieval-augmented schemes: Pairing a suspicious sample (text, image, audio) with a large database of known AI or human samples can reveal nearest neighbors, significantly enhancing detection of paraphrased or style-shifted outputs [21, 64]. (2) Cross-modal verification: Advanced generative systems produce text, images, and audio in tandem (e.g., a news article plus a purported video). Checking cross-modal consistency (lip sync, textual references to images, etc.) can reveal mismatches [54]. (3) Human-in-the-loop collaboration: No fully automated solution is perfect. Incorporating expert or crowd-sourced human checks for borderline cases, along with user feedback loops, is likely essential in high-consequence scenarios (e.g., journalism, law enforcement) [47]. (4) Private and federated learning approaches: As privacy restrictions tighten, training robust detectors on distributed user data may rely on federated or privacy-preserving methods that do not require centralizing sensitive content [10]. (5) Neurosymbolic and knowledge-enhanced detection: Fusing symbolic reasoning or domain knowledge with deep neural detection may help catch subtle anomalies overlooked by purely data-driven methods. This might be particularly relevant for domain-specific content such as medical, legal, or scientific articles.

### v2

@@ I Introduction
Contributions and Unifying Taxonomy
What is new. Beyond synthesizing prior work, we contribute:
  • A cross-modal taxonomy unifying textual, visual, and audio detectors into five families: (i) signal/statistical (stylometry, pixel/spectral cues), (ii) model-likelihood (perplexity/curvature), (iii) supervised classifiers (fine-tuned detectors), (iv) provenance mechanisms (watermarking/fingerprinting and content credentials), and (v) retrieval- and consistency-based (nearest-neighbor retrieval, cross-modal agreement checks).
  • A threat model matrix mapping attacker capabilities to defenses across modalities (Table I), clarifying realistic deployment choices and gaps.
  • Actionable deployment notes for retrieval-based text detection (mitigating paraphrasing), temporal/physiological video forensics, and real-time multilingual audio detection.
  • Reproducible resources linking datasets, leaderboards, and open-source implementations (Appendix A).
@@ II-B Threat Models and Malicious Use Cases
TABLE I: Attacker capabilities vs. defenses across modalities (text / image-video / audio). ✓ likely effective; $\triangle$ partially effective or context-dependent; ✗ ineffective. “Prov.” includes watermarking/fingerprinting and provenance standards (e.g., C2PA).
Attacker capability Text defenses Image/Video defenses Audio defenses
Stat./LLM Sup./Ensemble Prov. Pixel/Freq. Temporal Prov. Pipeline/NN End-to-end Prov.
Zero-effort (raw model output)
Paraphrasing / style transfer $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$
Heavy compression / resampling $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$
Cropping / rotation / time-stretch $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$
Model-switching (unmarked models) $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$
Adversarial perturbations $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$
Cross-lingual / domain shift $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$
Splicing / synthetic-real mixing $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$
Several works show that paraphrasing can substantially degrade watermark detection, though more robust schemes continue to be proposed [28, 32, 23, 31, 75, 52]. Provenance standards like C2PA provide complementary, opt-in authenticity metadata [14].
@@ III-B1 LLM prompting and zero-shot methods
One popular approach leverages an external LLM to classify whether a piece of text is AI-generated or not [7]. With carefully crafted prompts, these zero-shot methods can achieve moderate accuracy quickly. However, sensitivity to prompt design and the adversarial gap between the LLM detector and the text generator are common issues [61]. Recent zero-shot probabilistic curvature methods (e.g., DetectGPT and Fast-DetectGPT) trade accuracy and compute in different ways [44, 6].
@@ III-B4 Watermarking for AI text
Cooperative watermarking modifies text generation at token selection time, embedding an imperceptible pattern in the distribution of words or punctuation [28, 56]. A verifier can detect such patterns after the fact. While promising for major industrial LLMs that adopt the standard, watermarking fails if malicious or open-source models do not embed it, or if paraphrasing disrupts the signal [23, 32, 52, 75]. Complementary provenance standards (e.g., C2PA) aim to capture capture-chain information when available [14].
@@ III-B5 Ensemble and multi-feature systems
III-B6 Retrieval-augmented defenses against paraphrasing (deployment notes)
[…]
@@ IV-B4 Watermarking for visual media
When the generative pipeline is compliant, watermarks are embedded. These vary from invisible spatial pixel encodings to frequency manipulations [28, 40]. However, simple image transformations, such as crop or rotation, can weaken naive watermarks [23, 59] unless specifically designed for robustness. Content credentials such as C2PA can attach tamper-evident provenance metadata on capture and edit chains [14].
IV-B5 Temporal/physiological forensics for video
Beyond per-frame artifacts, temporal and bio-signal cues are highly informative. Methods exploit phoneme-viseme inconsistencies in lip motion [2], remote photoplethysmography (rPPG) from subtle skin color changes [13], eye-blink and micro-expression patterns [38], optical flow coherence and head-pose/motion geometry [46, 5, 11, 65]. These cues are often more resilient to spatial filtering but can be sensitive to heavy compression and low frame rates.
@@ IV-C Datasets and Benchmarks
Well-known datasets include CelebA-HQ [27] for facial images and LAION [58] for broad image domains. Large-scale synthetic benchmarks like GenImage (1M+ pairs; diverse generators and degradations) offer cross-generator and degraded-image evaluation [76, 77]. Community-maintained leaderboards and consolidated testbeds (e.g., AIGCDetect Benchmark) enable reproducible comparisons and track generalization to newer models [18, 20]. For video, FaceForensics++ [54], DFDC [17], and Celeb-DF are commonly used [64]. However, frequent advancement of the generative model requires continuous expansion of the detection dataset.
@@ V-B1 Pipeline classifiers
Features such as Mel frequency cepstral coefficients (MFCC), Linear frequency cepstral coefficients (LFCC) spectrogram-based descriptors are extracted and then fed into machine learning models (SVM, XGBoost, or CNN) [73, 34]. They typically rely on analyzing subtle cues in pitch, timbre, and vocal fold dynamics.
@@ V-B4 Audio watermarking
V-C Operational considerations: latency and multilingual robustness
Streaming/real-time: For telephony fraud prevention and live moderation, detectors must operate with very low latency. Sub-second chunking (e.g., 0.5–1.0 s hops with 2–3 s context windows) enables timely alerts, while causal CNN or Conformer backbones with look-ahead $\leq$200 ms help maintain responsiveness. Incorporating on-device VAD (voice activity detection) further reduces computational overhead and lowers false-alarm rates. Calibration: Detector performance should be tuned to in-domain EERs (equal error rates). Dynamic thresholding strategies that adapt to input SNR levels can significantly improve stability in noisy or variable conditions. Multilingual/broad-accent: Results from the ASVspoof evaluations, which include the Logical Access (LA) track for algorithmically generated speech, the Physical Access (PA) track for replayed audio, and the Deepfake (DF) track for highly realistic neural synthesis, consistently show sharp performance drops under cross-corpus and cross-condition settings. These findings underscore the importance of training with diverse languages and accents, followed by periodic domain adaptation [67, 70, 3]. Robustness can be further enhanced with self-supervised pretraining methods (e.g., HuBERT, Wav2Vec2) and test-time augmentations such as noise injection or codec simulation.
@@ VII-A Unified Themes Across Modalities
Despite modality differences, several consistent themes arise.
  1. 1.
    Arms race with generators: As generative models advance, older detection strategies degrade, requiring frequent retraining or adaptation [62, 43].
  2. 2.
    Vulnerability to perturbations: Slight paraphrasing in text, minor image edits (e.g., crop and color shift), or audio pitch/time adjustments can bypass naive methods [30, 55].
  3. 3.
    Watermarking as partial solution: Watermarking is useful if widely adopted by model providers, but it’s easily circumvented by uncooperative or malicious providers [32, 52].
  4. 4.
    Ensemble or multimodal methods: Combining multiple cues (statistical, watermark, contextual) and bridging text–image–audio modalities yields more robust detection [48, 74].
  5. 5.
    Ethical pitfalls: Systemic large-scale scanning can infringe on user privacy, and erroneous misclassifications can harm reputations.
VII-B Deeper analysis of watermark robustness
Empirical and theoretical studies report mixed results on watermark robustness under paraphrase, editing, and translation. While Unigram/semantic variants and adaptive schemes seek provable or empirical robustness [75, 32, 23], other analyses demonstrate learnability or reverse-engineering of green-list partitions and sharp drops in detection rates under targeted paraphrasing [52, 29]. We therefore position watermarking as one ingredient in a broader provenance strategy alongside content credentials (C2PA) and out-of-band logging for cooperative platforms [14].
VII-C Concrete pathways to address open challenges
Adversarial robustness: Strengthening resilience requires transformation-consistent training (such as paraphrasing or back-translation for text, crop/resize/codec chains for images, and time-stretching or noise perturbations for audio) combined with adversarial example mining and confidence calibration through temperature scaling. Domain generalization: To cope with unseen generators and distribution shifts, promising strategies include generator-aware data augmentation, continual learning with replay of past samples, and retrieval-based regularizers that leverage nearest-neighbor consistency. Explainability: Enhancing interpretability involves highlighting salient tokens, regions, or frames and linking them to human-understandable cues, for example, viseme–phoneme alignment [2] in video or rPPG signals in facial imagery. Governance: Beyond algorithmic techniques, standardized disclosures and verifiable audit trails, implemented via provenance metadata (e.g., C2PA manifests), should be paired with opt-in watermarking whenever feasible to ensure accountability and transparency.
VII-D Dataset gaps and mitigation
Current benchmarks are heavily skewed toward faces and English text, leaving major gaps in other domains. Non-face imagery (e.g., documents, user interfaces, medical scans, remote-sensing data) and low-resource languages remain underrepresented, limiting generalization to diverse real-world scenarios. In audio, multi-lingual, multi-accent, and codec-diverse corpora are still scarce despite notable progress from ASVspoof [67, 70]. To close these gaps, we recommend three complementary strategies: (i) targeted data collection in underrepresented modalities and languages; (ii) synthetic hard-negative generation to expose detectors to adversarially challenging examples; and (iii) cross-temporal benchmarks that explicitly test robustness against newly released generators, such as GenImage and evolving AIGCDetect testbeds [76, 18, 77].
@@ VIII Conclusion and Broader Reflections
Appendix A Reproducibility Resources (Datasets, Detectors, Scripts)
To support reproducibility and practical experimentation, we summarize widely used datasets, detectors, and code resources across modalities in Table II. This table centralizes pointers that can serve as entry points for researchers and practitioners.
TABLE II: Key reproducibility resources for AI-generated content detection.
Modality Representative Resources
Text DetectGPT, Fast-DetectGPT (zero-shot curvature) [44, 6]; retrieval-based defenses [30]; semantic/robust watermarks [23, 32, 75]
Image
/Video
FaceForensics++ [54]; DFDC [17]; GenImage (1M+ pairs) [76, 77]; community leaderboards [18, 20]; temporal/physiological cues such as viseme mismatch [2], rPPG [13], blink patterns [38]
Audio ASVspoof 2019/2021 datasets and challenge reports [67, 70, 3]; DeepMind SynthID audio watermark [19]
Provenance
/Policy
C2PA technical specification and overview [14]
Acknowledgments
We thank the reviewers of the MMAI Workshop on Multimodal AI, held in conjunction with the IEEE International Conference on Data Mining (ICDM) 2025, for their constructive feedback. Their comments directly motivated the addition of a unified taxonomy, a threat matrix, temporal-forensics notes, retrieval deployment guidance, and a consolidated resource appendix.